Foxx Company incurs $510000 overhead costs each year in its three main departments, setup ($20000), machining ($355000), and packing ($135000). The setup department performs 40 setups per year, the machining department works 5000 hours per year, and the packing department packs 500 orders per year. Information about Foxx’s two products is as follows: Product A1 Product B1 Number of setups 20 20 Machining hours 1000 4000 Orders packed 150 350 Number of products manufactured 600 400 If machining hours are used as a base under traditional casting, how much overhead is assigned to Product A1 each year? $255000 $177500 $153000 $102000
Answer #1

Total overheads = 20000+355,000+135,000 = 510,000

Total Machine Hours = 1000+4000 = 5000

Predetermined Overhead rate = Total overheads / total Machine hours = $510,000/5000 = 102

Overheads Applies to A1 = 102*1000 = 102,000.